# DreamFrame: Enhancing Video Understanding via Automatically Generated QA and Style-Consistent Keyframes

### v4

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–October 31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia(MM ’25), October 27–October 31, 2025, Dublin, Irelandisbn: 978-1-4503-XXXX-X/25/10ccs: Computing methodologiesccs: Computing methodologies Computer visionccs: Computing methodologies 2D Generationccs: Computing methodologies Video Understanding
@@ 1. Introduction
  • Leveraging our generative approach, we have developed and will publicly release a comprehensive dataset for video understanding, alongside a superb LVLM, i.e., DreamFrame-7B, trained for enhanced understanding. understanding

### v5

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–October 31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia($M$ ’25), October 27–October 31, 2025, Dublin, Irelandisbn: 978-1-4503-XXXX-X/25/10ccs: Computing methodologiesccs: Computing methodologies Computer visionccs: Computing methodologies 2D Generationccs: Computing methodologies Video Understanding
@@ 1. Introduction
  • Leveraging our generative approach, we have developed and will publicly release a comprehensive dataset for video understanding, alongside a superb LVLM, i.e., DreamFrame-7B, trained for enhanced video understanding.